Tech Staff Meeting Notes October 19, 2004

Current Top Priorities are:I-40 Bridge, BB Upgrade, TZTN


Greg's list

Routine Maintenance. DBB PM's. Every station serviced. DBB PM primarily complete. Done for 04.

BB network upgrade. VHF transceivers failed and have to be redesigned and rebuilt. Should have shipped 10/15. Need to work out telem paths.

Battery charging station. Need a hood. What to do about a body wash?

All strongmo stations serviced within past 6 months. On regular rotation.


Need to schedule trip to help Wurth install 3 stations.
Need to get malfunctioning systems back in order to install.
Need to revisit getting continuous connections to all possible stations.
Two stations down, John F is looking into it.

I-40 Bridge. Tentative installation to begin 10/21. All 7 recorders functioning together with mseed creation and xfer success. Greg to send detailed list upgrades to AZTN to Mitch to secure permission. Brewer will order DSL for AZTN after Mitch secures upgrade permission.

MKTA AC. Will replace. Done.

FDL. MSH deployment running smoothly. John F verified performance of remaining instruments to assure readiness for local deployments. Horton will get with John to help organize things to make sure we are ready to go.

Second Tier Priorities

FDL flooding. P-plant is allegedly on this.

CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. Filipcic needs to install.

MPHB. Mitch needs to determine current orientation so we can pull it and orient correctly.

BB Test Facility. Need to pour the slabs, install the conduit and electrical enclosure, etc. Now that soil voids are taken care of. Now what to do about the voids between the bottom of the vault and the concrete? Should we rent a tool to "settle" the ground around the vault (not this project does not have a budget). Done. Need to install wiring and electronics.

VBMS. Proposal should be forthcoming in a couple weeks.

SEKY (now TZTN). Power installed. Installation costs billed to CERI, monthly utility bill to USGS DFC. Need to get hardware from Golden and schedule installation.

Timing. Greg has an easy way to generate a datastream that goes high when lock is lost. Need a worm module to monitor. BB monitoring is more difficult. How do we automate the status stream or even get it back to Memphis for logging? (Save it to file with Scream?). Note, new version of firmware sends status to both flash and stream. Hence, information will be in gcf2ew logs, but how do we monitor that? Maybe not a huge deal with broadbands since both Jim and Steve regularly monitor. Would still be good to have a record though.

Cowcamps pit is done. A tower is being constructed at the top of the mountain that have permission to get on.

Blohole property now in the hands of a private landowner. Will allow continued operation of the station but no vehicles (i.e. a 1.5 mile hike). John as some leads on an alternative site.


North Carlina station refurbish. Keep node at LRNC. Modify to be the same as a New Madrid node and minimize internet bandwidth required.

Calibrate S-13's.

Nagios. Reno uses this to monitor SOH for a variety of systems including nodes and seismic stations. Could be very useful.

Gural EDU seismometers received. Works fine. Greg is working on power systems. Done?

Giddings tune-up.

Spare keys for new pickup.